OpenStudy (anonymous):

Solve. Please show all of your work. The second angle of a triangle is 4 times as large as the first. The third angle is more than the sum of the other two angles. Find the measure of all three angles.

OpenStudy (bahrom7893):

x+4x+y = 180

OpenStudy (bahrom7893):

hmm i think it's trial and error now.. x = 10 y = 40 x+y = 50 third angle z is probably 130

OpenStudy (bahrom7893):

i mean there are many options for this.

OpenStudy (bahrom7893):

it just said greater than, unless u wrote the question wrong.

OpenStudy (anonymous):

sorry!!! the question is: Solve. Please show all of your work. The second angle of a triangle is 4 times as large as the first. The third angle is 80 degrees more than the sum of the other two angles. Find the measure of all three angles.

OpenStudy (bahrom7893):

x+4x+(x+4x+80) = 180 5x + 5x + 80 = 180 10x = 100 x = 10 y = 40 z = 130 hmm i did get it right by guesswork.

OpenStudy (anonymous):

yes you did!!! cool :)

OpenStudy (anonymous):

f + s + t = 180 f + 4f + (5f + 80) = 180 10f = 180 - 80 -->f = 100/ 10 = 10 -> s = 4 * 10 = 40 => t = 5* 10 + 80 = 130 Pls double check your post ALWAYS!